40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 10 19:03:57 2011, rana, Update, LSC, Y Arm work 
    Reply  Sat Sep 10 19:45:29 2011, Jenne, Update, LSC, Y Arm work 
    Reply  Tue Sep 13 16:57:44 2011, steve, Update, LSC, ITMX oplev ** 
Message ID: 5382     Entry time: Sat Sep 10 19:45:29 2011     In reply to: 5381
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Y Arm work 

Quote:

 * ITMY OL: Also the optical layout seems strange: there are two copies of the laser beam going into the chamber (??).

* The ETMY OL beam was coming out but clipping on the mount for the ETMY OL HeNe. This indicates a failure on our part to do the ETMY closeout alignment properly.

 The 2nd beam from this laser is for the SRM's OpLev, so that shouldn't be changed.

For better or worse, we didn't do anything to the ETM OpLevs, because they don't have any in-vac steering optics.  We did however go through and check on all the corner OpLevs.

ELOG V3.1.3-